it took a motorboat 2 h more to travel 32 mi against a current than it took to travel 32 mi with the current. the rate of the current was 4 mph. find the rate of the boat in calm water
Answer by mananth(16946)

spee with wind x+4 mph
speed against wind=x-4 mph
Total flying time = 5.33 hours
Time with wind= -32/(x+4 )
time against wind 32/(x-4)
time against wind - time with wind = 2
-32/(x+4)+32/(x -4)=2
LCD =(x+ 4)*(x-4)
multiply the equation by the LCD
we get
-32*(x-4 )+32(x+4)=2(x^2-16)
-32x128+ 32x+128 )=2X^2-32
0x+256=2 x^2-32
2x^2-256 -32=0
2x^2-288 =0
/2
X^2-144 =0
x^2=144
x= +/- 12
x = 12 mph the boat speed